# Is Causality Necessary for Efficient Portfolios? A Computational Perspective on Predictive Validity and Model Misspecification

### v3

@@ 4.4 Sensitivity to Alignment
(a) Frontiers contract as $\rho$ decreases.
(b) Sharpe ratio scales linearly with $\rho$.
Figure 5: Sensitivity to Alignment. (a) Efficient frontiers shrink smoothly as alignment $\rho$ declines. (b) Sharpe ratios scale linearly with $\rho$, confirming the theoretical prediction.
@@ 4.5 Empirical Validation on Equity Data
(a) Mean–volatility scatter.
(b) Empirical frontier.
Figure 6: Real-World Signals from S&P 500 Stocks. (a) Scatter of assets by mean and volatility. (b) Empirical efficient frontier remains smooth and convex, showing that directional signals yield coherent portfolios even under misspecification.
@@ 4.6 Monte Carlo Validation of Bias Attenuation
Varying the confounding noise parameter $\sigma_{\zeta}$, we regress $Y=\beta X+\gamma\widetilde{Z}+\varepsilon$ on $X$ alone. Figure 7 shows empirical OLS estimates align with the theoretical attenuation curve from Equation (1). This confirms Corollary 1: omitted-variable bias degrades signals smoothly, not catastrophically.
(a) Monte Carlo estimates vs. theory.
$\sigma_{\zeta}$ $\hat{\beta}_{\mathrm{OLS}}$ $\beta_{\mathrm{theory}}$
@@
(b) Selected numerical values.
Figure 7: Attenuation of Omitted-Variable Bias. (a) Monte Carlo simulations (dots) align with the theoretical curve (dashed). (b) Numerical estimates match analytical predictions across noise levels, showing that omitted-variable bias diminishes smoothly as $\sigma_{\zeta}$ increases.
@@ 5 Discussion
This paper has revisited the claim that causal factor models are a prerequisite for portfolio optimization. Through theoretical results, simulation counterexamples, and empirical validation, we show that this assertion does not hold in general. Even when structurally misspecified, predictive models can produce signals that remain directionally aligned with true returns and sufficient to sustain smooth, convex efficient frontiers.
Our analysis clarifies several distinctions often blurred in recent debates. First, we separate directionality, ranking, and calibration. Theorem 1, Lemma 2, and Corollary 2 establish that directional correctness and ranking are necessary but not sufficient for efficiency: miscalibration attenuates realized returns and Sharpe ratios. This explains why predictive accuracy in classification does not automatically translate into optimization efficiency, echoing the explanatory–predictive divide highlighted by Shmueli, (2010).
Second, we refine the interplay between ranking and calibration. In the sort-based framework of Almgren and Chriss, (2005), correct ranking ensures Pareto efficiency. Proposition 1 and Lemma 3 extend this by showing that calibration governs quantitative performance within the efficient set. Figures 2 and 5 confirm empirically that while ranking secures validity, calibration determines realized efficiency.
Third, we counter the claim that misspecification necessarily collapses optimization geometry. Prior critiques (e.g. (López de Prado et al.,, 2024, 2025)) argued that omitted variables or nonlinearities produce inverted signals and degenerate frontiers. Our counterexamples in Sections 3.2 and 3.4 show the opposite: predictive signals often retain directional informativeness, and frontiers remain convex and non-degenerate.
@@
Fourth, Section 3.5 demonstrates that signal viability is not knife-edge. Both analytically and empirically, omitted-variable bias attenuates continuously as dependence weakens (Figure 7). This continuity result highlights robustness: efficiency degrades smoothly rather than collapsing.
@@
Finally, our empirical study on S&P 500 equities (Figure 6) shows that predictive signals yield coherent frontiers in practice. Even without structural correctness, directional validity with approximate calibration suffices for actionable portfolio design. Predictive modeling should therefore be viewed as a complementary and often more pragmatic tool when structural assumptions are uncertain, unstable, or untestable.
While causal inference remains valuable for attribution, regime analysis, and risk decomposition (Lo,, 2017; Wilcox and Gebbie,, 2014; Rodriguez Dominguez,, 2023), our results demonstrate that causal identifiability is not a prerequisite for optimization efficiency. The decisive benchmarks are whether signals preserve directional informativeness, maintain correct ranking, and achieve sufficient calibration. By formalizing this decomposition and validating it empirically, we provide a constructive correction to structural critiques and advocate for evaluation criteria that recognize the robustness and operational utility of predictive-first approaches.
@@ 6 Conclusion and Future Work
We have offered a systematic re-evaluation of the claim that causal factor models are necessary for portfolio efficiency. Through formal analysis, controlled simulations, and empirical evidence, we show that structural misspecification, such as omitted variables or nonlinear confounding, does not inherently invalidate optimization. Predictive models without causal identifiability can nevertheless yield directionally valid signals that sustain smooth and convex efficient frontiers.
@@
Our contributions are fourfold. First, we disentangle directionality, ranking, and calibration, proving that directional correctness is necessary but insufficient when magnitudes are miscalibrated. Second, we show that structurally misspecified yet directionally aligned signals preserve optimization geometry, extending the portfolios-from-sorts framework by Almgren and Chriss, (2005) to account for calibration effects. Third, we demonstrate analytically and empirically that signal viability is robust, with omitted-variable bias attenuating smoothly rather than collapsing abruptly. Fourth, we validate these findings with real market data, confirming that predictive signals yield coherent frontiers even absent strict causal grounding.
These results caution against overstating the necessity of causal inference in portfolio design. While causal models remain indispensable for attribution, risk decomposition, and policy analysis (Lo,, 2017; Wilcox and Gebbie,, 2014; Rodriguez Dominguez,, 2023), their superiority is not automatic. What ultimately governs optimization performance is the directional fidelity and calibration of predictive signals, not structural correctness per se. This perspective motivates hybrid approaches that integrate causal diagnostics with predictive-first models to balance interpretability and robustness.
Future work should pursue three directions: integrating causal diagnostics directly into predictive pipelines, systematically benchmarking causal versus predictive models under regime shifts, and extending misspecification-resilient optimization theory to high-dimensional and nonlinear contexts. Such research can clarify when a causal structure enhances decision-making and when predictive validity alone is sufficient.
To sum up, structural correctness may be desirable, but it is not a prerequisite for portfolio efficiency. Effective portfolio construction relies on signals that are directionally valid, properly calibrated, and resilient to the inevitable imperfections inherent in financial modeling.
@@ Appendix A Simulation Protocols and Code
All simulations were implemented in Python using NumPy, Matplotlib, and scikit-learn where necessary. Random seeds were fixed for reproducibility. Across experiments:
@@ Appendix B Refutation of Structural Necessity Claims
Several recent critiques of predictive approaches to portfolio optimization (López de Prado et al.,, 2024, 2025) have advanced strong structural arguments suggesting that causal factor models are not only desirable but strictly necessary for efficiency. These arguments rest on claims such as: (i) misspecification inevitably produces sign inversion; (ii) sign inversion is the sole driver of inefficiency; (iii) sign agreement guarantees efficiency; and (iv) coherent efficient frontiers cannot exist without fully specified causal models.
@@
Misspecification necessarily causes sign inversion $\text{Misspecification}\Rightarrow\text{Sign Inversion}$ Counterexamples in Sections 3.2 and 3.4 show that alignment can persist under omitted-variable bias. Figures 1 and 3 confirm empirically that inversion is not inevitable.
@@
Misspecification invalidates mean–variance optimization $\text{Misspecification}\Rightarrow\text{No Frontier}$ Lemma 4 and Corollary 3 prove that positive alignment suffices for smooth, convex frontiers. Section 4 (Figures 4, 5) confirms this empirically.
@@
Table 2: Structural claims from (López de Prado et al.,, 2024, 2025) and their refutations. Across cases, inefficiency is shown to arise from miscalibration rather than inevitable inversion, and viable frontiers persist under positive alignment.
@@ Appendix C Correction of Variance Normalization in Factor Models
In Section 3.1, we identified a normalization discrepancy in (López de Prado et al.,, 2024), where the reported variance of $F_{2}$ is inconsistent with standard factor modeling assumptions. We now provide the corrected derivation and its implications for bias interpretation.
@@ C.2 Comparison
(López de Prado et al.,, 2024): $\mathrm{Var}(F_{2})=1+\delta^{2}$ $\hat{\beta}_{n}=\dfrac{\beta_{n}+\gamma_{n}\delta}{1+\delta^{2}}$ Nonlinear attenuation; no longer proportionate to confounding strength

### v4

@@ 4.4 Sensitivity to Alignment
(a) Efficient frontiers constructed from surrogate signals with decreasing cosine alignment $\rho$ relative to true expected returns. As alignment weakens, attainable risk–return trade-offs contract smoothly without loss of convexity or feasibility.
(b) Population Sharpe ratios as a function of cosine alignment $\rho$. Efficiency declines linearly with misalignment, confirming the theoretical scaling law derived in Proposition 2.
Figure 6: Sensitivity of Portfolio Efficiency to Directional Alignment. Panel (a) shows that reduced alignment leads to a proportional contraction of the efficient frontier while preserving its geometry. Panel (b) shows that Sharpe ratios scale linearly with alignment. Together, the panels demonstrate that efficiency loss under misspecification arises through smooth geometric contraction rather than frontier collapse.
@@ 4.5 Empirical Validation on Equity Data
(a) Cross-section of S&P 500 assets. Each point represents a stock, plotted by realized mean return (y-axis) and volatility (x-axis). The dispersion reflects heterogeneous risk–return profiles and provides the empirical geometry on which portfolio diversification operates.
(b) Empirical efficient frontier under predictive signals. Portfolios constructed from predictive (potentially misspecified) signals trace a smooth, convex risk–return frontier. This demonstrates that directional alignment is sufficient to preserve optimization geometry in real data.
Figure 7: Empirical validation using S&P 500 equities. Panel (a) shows the realized cross-sectional risk–return distribution of assets. Panel (b) shows the corresponding empirical efficient frontier obtained via mean–variance optimization using predictive signals. Despite potential misspecification and lack of causal identification, the frontier remains well defined and convex, confirming that directional validity and approximate calibration suffice for actionable portfolio construction in practice.
@@ 4.6 Monte Carlo Validation of Bias Attenuation
Varying the confounding noise parameter $\sigma_{\zeta}$, we regress $Y=\beta X+\gamma\widetilde{Z}+\varepsilon$ on $X$ alone. Figure 8 shows that empirical OLS estimates align closely with the theoretical attenuation curve from Equation (1), while Table 2 provides numerical confirmation across noise levels. Together, these results confirm Corollary 1: omitted-variable bias degrades predictive signals smoothly through attenuation rather than inducing catastrophic sign inversion.
To complement the empirical analysis, Appendix E presents high-dimensional simulation-based stress tests under controlled misspecification, designed to isolate the roles of directional alignment, calibration, and constraints independently of causal structure.
Figure 8: Continuous attenuation of omitted-variable bias under confounding noise. The horizontal axis reports the standard deviation of idiosyncratic confounding noise $\sigma_{\zeta}$. Dots represent Monte Carlo estimates of the misspecified OLS coefficient $\widehat{\beta}_{\mathrm{OLS}}$, while the dashed line shows the theoretical attenuation curve from Equation (1). Increasing confounding noise attenuates the coefficient smoothly toward the true value, illustrating that omitted-variable bias weakens continuously rather than inducing sign inversion.
Table 2: Numerical verification of attenuation law. Monte Carlo estimates of the misspecified OLS coefficient closely match analytical predictions across confounding-noise levels, confirming the exact bias-decay formula derived in Proposition 3.
$\sigma_{\zeta}$ $\widehat{\beta}_{\mathrm{OLS}}$ $\beta_{\mathrm{theory}}$
@@ 5 Discussion
This paper revisits the claim that causal factor models are a prerequisite for portfolio optimization. Through theoretical results, simulation counterexamples, and empirical validation, we show that this assertion does not hold within the static mean–variance and closely related quadratic portfolio optimization frameworks analyzed in this paper. Even when structurally misspecified, predictive models can produce signals that remain directionally aligned with true returns and are sufficient to sustain smooth, convex efficient frontiers.
5.1 On the Claim That Misspecification Necessarily Implies Signal Inversion
A recurring assertion in recent critiques is that structural misspecification, including omitted variables and nonlinear confounding, inevitably produces inverted predictive signals and collapses the efficient frontier. Our results directly contradict this claim within the static mean–variance portfolio optimization framework analyzed in this paper. Sections 3.2 and 3.4 provide explicit counterexamples showing that omitted-variable bias can induce attenuation without inversion, preserving directional informativeness under moderate departures from the true data-generating process.
@@ 5.1 On the Claim That Misspecification Necessarily Implies Signal Inversion
Section 3.5 further demonstrates that signal degradation under misspecification is not knife-edge. Both analytically and empirically, omitted-variable bias attenuates continuously as dependence weakens (Figure 8 and Table 2). This continuity result highlights a precise notion of robustness: optimization performance degrades smoothly rather than collapsing abruptly.
5.2 On the Claim That Inefficiency Requires Signal Inversion
A second, closely related claim is that inefficiency in portfolio optimization arises only through sign inversion. Our analysis clarifies why this is incorrect within the static mean–variance framework. We explicitly separate directionality, ranking, and calibration. Theorem 1, Lemma 2, and Corollary 2 establish that directional correctness and ranking are necessary but not sufficient for efficiency: miscalibration alone is sufficient to attenuate realized returns and Sharpe ratios.
This refines the interpretation of the sort-based framework of Almgren and Chriss (2005). While correct ranking ensures efficient-set membership, Proposition 1 and Lemma 3 show that calibration governs quantitative performance within that set. Figures 3 and 6 confirm empirically that inefficiency most often arises through magnitude distortion rather than sign failure.
5.3 On the Robustness of Optimization Geometry
Beyond signal properties, we examine whether optimization geometry itself is fragile to misspecification. Lemma 4 and Proposition 2 show that, as long as surrogate signals remain directionally aligned with true expected returns, the efficient frontier remains smooth, convex, and non-degenerate within the static mean–variance optimization problem with a covariance risk model. Structural misspecification does not collapse optimization geometry.
In this context, robustness refers to the preservation of the mean–variance efficient frontier geometry under misspecified inputs, rather than invariance of optimal weights or realized returns. The static mean–variance problem remains well posed under a wide class of misspecifications: feasibility and convexity are preserved, while investment efficiency degrades smoothly through miscalibration rather than failing discontinuously.
@@ 5.3 On the Robustness of Optimization Geometry
Our empirical study on S&P 500 equities (Figure 7) confirms that these theoretical properties persist in practice. Even without structural correctness, directional validity combined with approximate calibration suffices for coherent and actionable portfolio design.
5.4 On the Role of Causality Beyond Optimization
The results of this paper do not argue against the use of causal inference in finance. Causal models remain essential for tasks such as structural attribution, regime analysis, stress testing, and counterfactual reasoning. The contribution of this work is instead to clarify a boundary condition: causal identifiability is not a prerequisite for portfolio efficiency within the static mean–variance and closely related quadratic portfolio optimization frameworks, provided that predictive signals satisfy empirically verifiable geometric conditions.
From an operational perspective, the relevant object for portfolio construction is not causal structure itself, but the empirical geometry induced by predictive signals in interaction with the risk model. Portfolio optimization remains viable whenever surrogate return vectors preserve positive directional alignment and stable ranking relative to realized returns. Structural misspecification primarily affects performance through miscalibration of signal magnitudes rather than systematic sign inversion.
This distinction has direct practical consequences. Predictive models (whether causally motivated or purely associational) can be evaluated for portfolio suitability using observable diagnostics, including out-of-sample directional alignment, ranking stability across rolling windows, and sensitivity of portfolio performance to signal rescaling. When these conditions are satisfied, optimization remains well posed, and calibration determines the attainable position along the efficient frontier.
Causal information becomes relevant for portfolio optimization only insofar as it informs the construction or stabilization of the diversification geometry itself, rather than as a requirement imposed on return prediction. Within any such geometry, investment efficiency continues to be governed by alignment, ranking, and calibration. This perspective provides a principled role for causality while preserving a clear separation between explanatory objectives and optimization feasibility.
@@ 6 Conclusion and Future Work
We have re-evaluated the claim that causal factor models are necessary for portfolio efficiency within the static mean–variance and closely related quadratic portfolio optimization frameworks. Through formal analysis, controlled simulations, and empirical evidence, we show that structural misspecification, including omitted variables and nonlinear confounding, does not inherently invalidate portfolio optimization. Predictive models without causal identifiability can nevertheless yield directionally informative signals that sustain smooth and convex efficient frontiers.
@@
Our contributions can be summarized as follows. First, we disentangle directionality, ranking, and calibration, establishing that directional correctness is necessary but not sufficient for portfolio efficiency when signal magnitudes are miscalibrated. Second, we show that structurally misspecified yet directionally aligned signals preserve optimization geometry, extending the portfolios-from-sorts framework of Almgren and Chriss (2005) by explicitly accounting for calibration effects within the efficient set. Third, we demonstrate analytically and empirically that misspecification induces smooth attenuation rather than abrupt collapse of efficiency. Fourth, we validate these results with real market data, confirming that predictive signals can support coherent and actionable frontiers even in the absence of strict causal grounding.
These findings caution against overstating the necessity of causal inference in portfolio design. While causal models remain indispensable for attribution, regime analysis, and stress testing (Lo, 2017; Wilcox and Gebbie, 2014), causal identifiability is not intrinsic to the objective of portfolio efficiency within the static mean–variance and closely related quadratic optimization frameworks analyzed here. Optimization performance is governed by the geometric relationship between predictive signals and realized returns, specifically directional alignment, ranking stability, and calibration, rather than by structural correctness per se. More broadly, the results clarify a recurring misconception in the literature: improvements in causal validity or predictive accuracy at the asset level do not, by themselves, determine investment efficiency. Portfolio optimization is fundamentally a diversification problem defined relative to a chosen risk–return geometry. For example, causal information can be incorporated into portfolio construction by shaping the diversification geometry itself, such as through sensitivity-based representations of common causal drivers (Rodriguez Dominguez, 2023). In such approaches, causality informs the structure of the risk–return space, while investment efficiency remains governed by the geometric properties of the resulting optimization problem.
When causality is incorporated into portfolio construction, it operates by shaping or stabilizing the diversification geometry itself rather than as a requirement imposed on return prediction. Even in such settings, investment efficiency remains governed by alignment, ranking, and calibration within the induced space. Future work may extend these results in several directions, including integrating causal diagnostics into predictive pipelines without conflating them with efficiency criteria, examining the geometric sufficiency conditions under alternative risk functionals and non-quadratic objectives, and studying misspecification-resilient optimization in higher-dimensional and nonlinear settings.
@@ Appendix A Simulation Protocols and Code
All simulations were implemented in Python using NumPy, Matplotlib, and scikit-learn where necessary. Random seeds were fixed for reproducibility. The protocols described in this appendix support the numerical results in Section 4 and the high-dimensional stress tests reported in Appendix E. Across experiments:
@@
Simulation return mapping. In simulation-based experiments that employ probabilistic classifiers (Sections 3.4 and 3.5), predictive models output class probabilities $p(x)=\mathbb{P}(Y=1\mid x)$, which are mapped to realized returns via the affine transformation $r(x)=a\,(2p(x)-1)+\varepsilon$, with $a>0$. This mapping preserves directional alignment, ranking, and cosine geometry with respect to the true signal, ensuring consistency with the linear mean–variance framework analyzed in the theory. All results in these experiments depend only on directional alignment and calibration, not on the specific affine scaling.
@@ A.4 Code for Bias Attenuation Simulation
    # sigma_eta^2 fixed to 1 and absorbed into baseline normalization
@@ Appendix B Refutation of Structural Necessity Claims
This appendix corresponds to the refutation roadmap outlined in the Introduction and summarized in Table 3, providing formal and empirical resolution of the necessity claims discussed in the main text within the class of static portfolio optimization frameworks considered in this paper.
Several recent critiques of predictive approaches to portfolio optimization (López de Prado et al., 2024, 2025a) have advanced strong structural arguments suggesting that causal factor models are not only desirable but strictly necessary for efficiency. These arguments rest on claims such as: (i) misspecification inevitably produces sign inversion; (ii) sign inversion is the sole driver of inefficiency; (iii) sign agreement guarantees efficiency; and (iv) coherent efficient frontiers cannot exist without fully specified causal models.
@@
Importantly, these conclusions do not assert that causal structure is irrelevant, but rather that causal identifiability is not a necessary condition for portfolio efficiency within the optimization settings analyzed here.
@@
Misspecification necessarily causes sign inversion $\text{Misspecification}\Rightarrow\text{Sign Inversion}$ Counterexamples in Sections 3.2 and 3.4 show that alignment can persist under omitted-variable bias. Figures 2 and 4 confirm empirically that inversion is not inevitable.
@@
Misspecification invalidates mean–variance optimization $\text{Misspecification}\Rightarrow\text{No Frontier}$ Lemma 4 and Corollary 3 prove that positive alignment suffices for smooth, convex frontiers. Section 4 (Figures 5, 6) confirms this empirically.
@@
Table 3: Structural claims from (López de Prado et al., 2024, 2025a) and their refutations. Across cases, inefficiency is shown to arise from miscalibration rather than inevitable inversion, and viable frontiers persist under positive alignment.
@@ Appendix C Correction of Variance Normalization in Factor Models
In Section 3.1, we identified a normalization discrepancy in (López de Prado et al., 2024), where the reported variance of $F_{2}$ is inconsistent with standard factor modeling assumptions. We now provide the corrected derivation and its implications for bias interpretation.
@@ C.2 Comparison
(López de Prado et al., 2024): $\mathrm{Var}(F_{2})=1+\delta^{2}$ $\hat{\beta}_{n}=\dfrac{\beta_{n}+\gamma_{n}\delta}{1+\delta^{2}}$ Nonlinear attenuation; no longer proportionate to confounding strength